# DistALANER: Distantly Supervised Active Learning Augmented Named Entity Recognition in the Open Source Software Ecosystem

### v3

@@ 6.1 Stage 1: Construction and matching of dictionary
Package (PKG)
pypy-configparser, account-plugin-twitter, gtkhtml3.2, pdfcrack, libfields-camlp4-dev-pf4q7
140062
Operating System (OS)
SymbOS, Unix System III, NOS, Windows, Cosmic, kubuntu
877
Organization (ORG)
launchpad, bugzilla, sourceforge, nokia, HP
379
Command (CMD)
alias, arch, bzip, cat, clear
141
Error (ERR)
No such process, No child processes, EFAULT, Bad address, EFBIG
124
Extension (EXT)
.asm, .gz, .html, .log, .php
76
Peripheral (PRP)
keyboard, mouse, printer, scanner, microphone
23
Software component (SOC)
bios, driver, ui, ntfs, fat32
12
Architecture (ARC)
x86, x64, 32-bit, 64-bit, amd64
7
@@
6.2 Stage 2: Entity distillation and dictionary expansion
In Stage 2, we distill the exactly matched entities and then expand the list of entities for a given entity type using an active learning approach. Active learning enables our system to learn iteratively, refining its understanding as it receives feedback from its interactions. We observe that many phrases get marked as an entity, but in reality, some of them are not entities. Thus, we employ two heuristics to distill the entities: (1) identifying parts-of-speech patterns, and (2) human intervention. Parts-of-speech patterns: After performing exact matching, we find that the likelihood of a phrase being an entity often depends on the parts-of-speech tag of the words in the phrase or before and after the phrase. For instance, the verb ‘find’ in a sentence is never an entity, even though our dictionary matching may mistakenly identify it as a CMD entity (see more samples in section 7). We adjust the contents of the dictionary through such iterative revisions. Human intervention: In this step, we avail human intervention to identify cases where a phrase has been incorrectly tagged as an entity. Active learning is particularly useful here, as the model can learn from the feedback provided by the human expert, improving its future predictions. We randomly sample a few automatically annotated bugs from all the bugs and check which phrases could potentially be non-entities. We encounter a challenge as the number of entities for categories like software components and peripherals is quite low in our entity list. To expand this dictionary, we utilize the software tool, TagMe [^16] https://sobigdata.d4science.org/web/tagme/tagme-help, to extract mentions from Wikipedia and consider them as entities. Along with that process, after receiving the annotated data from Stage 1 and Parts-of-speech filtered bugs (L) also with their IO mentions, a binary RoBERTa-based classifier is trained. We sample 100 bugs on a yearly basis, with our sample spanning from 2004 to 2019. Subsequently, we take up 100*16 = 1600 bugs from TagMe but not included in Stage 1 and feed them into our model for classification as either “entity” or “non-entity”. If the confidence for an entity being identified is 50% or greater, it undergoes manual labeling by a human to categorize it into one of the nine predefined entity types. Once the manual annotation is completed, the newly labeled data (d) is integrated into L, resulting in an updated dataset, (L $\cup$ d). The model is then retrained using this revised dataset, and the entire procedure is reiterated until the stopping criterion is met. The stopping criterion is defined as the absence of data in the list of entities identified by TagMe. Finally, our requirement for human intervention end up with around $\sim$3682 filtered mentions, a relatively small number considering the size of our dataset which contains 170K bugs (equivalent to $\sim$1.2 million entities). Distributions of identified entities are shown in Figure 3.
@@ 6.3 Stage 3: The NER model
We use a variety of NER models for obtaining the final tags. These include Linear-CRF Lafferty et al. (2001), BilSTM-CRF Huang et al. (2015), BERT-CRF Souza et al. (2020), BERT-NER Liu et al. (2021), RoBERTa-CRF Jurkiewicz et al. (2020), SpanBERT-CRF Portelli et al. (2021), and SoftNER Tabassum et al. (2020). All models are implemented to execute task-specific actions, with their performance evaluated using precision, recall, and F1-score. However, recall is highlighted in this paper as a representational metric for NER tasks. We train the existing CRF based models using our distantly labelled data obtained from Stage 1 and Stage 2. We note the different hyperparameters for the above models in Appendix D. We also use LLMs in a zero-shot setting with instruction based prompt to extract the entities along with its start and end index from the paragraph (see section 10 for more details).
@@ 7 Heuristics
When implementing NER using distant supervision, we use Part-of-Speech (POS) tagging as a heuristic to discard certain entities. In our process, we filter out the entities based on their POS tags. For example, we discard entities that are typically labeled as conjunctions, interjections, or prepositions as these are less likely to represent valid entities. This strategy ensures that only the most relevant words, such as nouns or proper nouns, are considered for entity recognition. This way, we can reduce noise in the data and improve the performance of our NER model. Furthermore, these POS tag-based heuristics help us in refining our entity list, leading to a more accurate and efficient distant supervision process in NER. In this section we enumerate some of the heuristics in Table 3 that we use to discard incorrect entities.
@@
"This results in a serious compromise on the possibility of running remote displays systems on ubuntu. In fact, the latter can only
@@
"SST will fail if donor has to send keyring. Looks like the donor is trying to send the file
while socat is still opening port $4$ on joiner: 20170606 09:00:15.294 WSREP_SST: [INFO] Streaming GTID
file before SST 20170606 09:00:18.368 WSREP_SST: [INFO] Streaming donor-keyring file before SST"
@@
8 Experimental setup
Training and test data preparation: In order to train the NER model in Stage 3, we need training data. For this purpose we select all the bug descriptions from the years 2004 through 2013 inspired by  Hazra et al. (2021, 2023a, 2023b). These descriptions, nearly 65K in number are automatically annotated using the first two stages of our framework. Thus, the training data only has auto-curated (aka silver) entity labels without any human involvement. For the test set we consider the data from the years 2016 to 2019. To assess the performance of the model, which has been trained using distantly labeled data, we perform human annotation of a subset of the test set.
Human annotation: We present 500 bug descriptions to four domain experts, each claiming over three years of experience in the Ubuntu ecosystem and package management. The resulting inter-annotator agreement among these annotators is 0.625 (see Appendix C for more information). For the Launchpad QA dataset, we employ the same group of experts to annotate the entities in 500 question-answer pairs, ensuring consistency in our data annotation process across both datasets.
Baselines: In addition to the NER models and zero-shot LLMs, we also use the combined output of Stage 1 and 2 as a baseline.
@@ 8 Experimental setup
Table 4: Recall rate for named entity recognition approaches. The first sub-row of each row shows the results for the Ubuntu bug dataset while the second sub-row shows the results for the Launchpad QA dataset (inference phase only). The best and the second best results in the first sub-row (bug dataset) are highlighted in dark and light green respectively. The best and the second best results in the second sub-row (QA dataset) are highlighted in dark and light purple respectively. BERT-CRF is significantly different from domain specific SoftNER (*$p=0.039$). Table 5 compares methods for HIn based on other metrics.
Metric for evaluation: To evaluate all the models, we compute the recall rate inspired by Tu and Lignos (2021). Recall rate is the proportion of actual entity types in ground truth that are accurately predicted. We use this metric rather than the F1 score in the main table in cognizance of the fact that human annotations are sometimes incomplete while the model is able to generate the correct entity type (see Section 9). Precisely, when dealing with software-related texts, overlooking an entity can lead to the omission of vital data. For instance, not recognizing an error code, a software package, or a particular function name can change the outcome from comprehending and addressing a problem to failing to do so. This makes capturing as many relevant entities as we can, which recall assesses, crucial. Further, software-related documents typically present crucial details that require exhaustive extraction for subsequent applications like bug tracking, requirement analysis, or code generation. Overlooking an entity in these contexts can have an adverse impact, underscoring the importance of recall as a key metric. Nevertheless, in order to make the results complete we also subsequently measure and report the precision, recall and F1 score.
9 Results
We evaluate all the baselines for both the Ubuntu bug dataset and the Launchpad QA dataset. The main results are noted in Table 4. Our experiments involve two different setups – (a) Human-Induced Training (HIn), and (b) Human-Only Training (HOn). In the case of a Human-Induced (HIn) configuration, we train the models using all the auto-annotated bug descriptions plus 10% (or 50) human-annotated bug descriptions. These examples of human annotated data are incorporated into the training to provide the models with additional knowledge of gold annotations. For Human-Only training (HOn), only 10% of human annotations are used to train the models (see subsection 9.2). We do not include any auto-annotated data during HOn configuration training. In both setups, 10% of the human-annotated data is used for training, 20% for validation, and 70% for testing.
@@ 9 Results
Methods Overall
Precision Recall Macro-F1
Linear-CRF 0.381 0.443 0.410
0.335 0.375 0.354
BiLSTM-CRF 0.268 0.314 0.290
0.077 0.111 0.090
BERT-NER 0.392 0.461 0.424
0.355 0.378 0.366
BERT-CRF 0.462 0.481 0.471
0.341 0.381 0.360
RoBERTa-CRF 0.443 0.452 0.448
0.325 0.361 0.342
spanBERT-CRF 0.314 0.396 0.350
0.298 0.367 0.330
SoftNER 0.377 0.416 0.396
0.296 0.346 0.319
Table 5: Comparison of methods for HIn based on precision, recall, and macro-F1 Scores. The first sub-row of each row shows the results for the Ubuntu bug dataset while the second sub-row shows the results for the Launchpad QA dataset (inference phase only). The best and the second best results in the first sub-row (bug dataset) are highlighted in dark and light green respectively. The best and the second best results in the second sub-row (QA dataset) are highlighted in dark and light purple respectively.
@@ 9.1 Progressive learning
Progressive learning Chatterjee et al. (2017); Moeed et al. (2020) is a strategy where machine learning models receive training in an incremental manner, with larger chunks of data over time. We conduct our progressive learning experiment by dividing our data into 25%, 50%, and 75% segments among equally distributed entities. Figure 4 presents the distribution of Recall rates across varying training data percentage. We limit this experiment to the top three performing models only, providing a focused perspective on how they react to this progressive learning approach. We observe that BERT-CRF continuously gains with additional data finally achieving the best performance with 100% training data.
@@ 9.2 Motivation of different split of HOn
Method ARC CMD ERR EXT OS ORG PKG PRP SOC Overall
BERT CRF (HOn) (Bug) 0.040 0.044 0.089 0 0.319 0.183 0.024 0.179 0.138 0.137
BERT CRF (HOn) (Launchpad) 0.037 0.043 0.042 0.063 0.194 0.174 0.032 0.090 0.134 0.089
Linear CRF (HOn) (Bug) 0.792 0.190 0.007 0.513 0.662 0.371 0.419 0.193 0.148 0.356
Linear CRF (HOn) (Launchpad) 0.088 0.140 0.028 0.218 0.181 0.027 0.029 0.009 0.019 0.068
Table 6: Comparison of performance metrics for BERT-CRF and Linear-CRF methods across various categories using different datasets (Bug and Launchpad).
@@ 10 Issues in finding distant labels using LLMs
In our current research, we employ a consistent instruction-based prompt to extract and tag entities from a given paragraph (see Table 9) for sample prompts). The goal is to return these tagged entities, along with their start and end indices, in a JSON format. The paragraph is analyzed to identify the nine entities. We evaluate the outputs from three different language models: GPT-3.5-Turbo, GPT-4, Google BARD and UniversalNER. Through our analysis, we make several observations: (a) GPT-4 fails to generate the complete JSON format, unlike GPT-3.5-Turbo and Google BARD; (b) The output pattern from GPT-4 and Google BARD exhibits more consistency than that of GPT-3.5-Turbo. For instance, when generating entities using GPT-4 and Google BARD, the model returns the entity, its type, and its start and end indices. However, GPT-3.5-Turbo occasionally returns the entity type and indices without the text itself, or merely refers to specific lines in the text leading to inconsistencies in its output structure; (c) In terms of content, both GPT-3.5-Turbo and GPT-4 extract more accurate entities compared to Google BARD. However, the indices produced by GPT-3.5-Turbo and GPT-4 tend to contain more errors compared to those generated by Google BARD; (d) We notice that all three models sometimes invent their own entity types, which they subsequently return in the output; (e) In the recently published UniversalNER Zhou et al. (2023), the model struggles to generate accurate entities, despite being trained on relevant data. (e) Finally, access to these model predictions are far more expensive compared to our distantly supervised method. See Tables 11 and 12 (from Appendix) for some examples.
@@ 11 Additional experiments for task-based evaluation
Relation extraction and NER are closely intertwined in many NLP tasks. They are often referred to as “sister" problems in the literature. Many studies, including Wang et al. (2022); Zhong and Chen , explore these two tasks jointly, highlighting their interdependence. While relation extraction identifies and classifies the relationships between recognized entities, it often serves as a supplementary evaluation, building upon the foundational results of NER models.
@@ 11.1 Relation extraction
dependency), etc.), (b) conflict – conflict relation refers to a situation where two or more software components, packages, or entities cannot coexist or function harmoniously due to incompatibilities, overlapping functionalities, or conflicting configurations (e.g. triplets include (cups, Printers, conflicts), (Keyboard, Input Method Editor (IME), conflicts), etc.), affected version – these indicate which specific versions of a software are affected by reported issues or bugs (e.g. triplets include (Flatbed Scanner, macOS Mojave, affected version), (Error Code 134, sudo dpkg --configure -a, affected version), etc.), cause and effect – these relations refers to cases where one entity (cause) triggers another entity (effect) (e.g. triplets include (Error Code 502, nginx, cause and effect), (Error Code 401, openssh-server, cause and effect), etc.), and interaction/control – corresponds to relations where an entity exerts dynamic influence on/manipulates another entity (e.g. triplets include (apt, install, interaction/control), (docker-ce, run, interaction/control), etc.).
Dataset: For this experiment we use the earlier 500 bug description data that was manually annotated for the named entities. This data is further annotated with the relationships among the entities by 3 expert annotators with an inter-annotator agreement of 0.693. We end up with a total of 642 triplets, each consisting of a head entity, relationship, and tail entity. Out of these, 27% have a dependency relationship, 21% are of type affected versions while conflicts, cause and effect, and interaction/control types account for 4%, 17%, and 31% of the triplets respectively. Since the data for conflict type is very less, we ignore it for our experiments.
@@
Results: In Table 7 we show the results of the relation extraction task which is posed as a classification problem having the triplet (the head entity, the tail entity and the relation type) representation as the input and one of the relation classes (dependency, affected versions, cause and effect, and interaction/control) as output. In the input, we draw the entity representations from our previously trained best NER models (i.e., trained BERT-CRF, trained BERT-NER, and trained RoBERTa-CRF) and compare their performance with vanilla BERT. Naturally, the rationale for employing the trained model as encoder is to acquire a more contextual representation than with the vanilla BERT. Further, we fine-tune the classifier layer with a very small amount of data (3%) to perform the relation classification. Dark cyan cells in Table 7 represent the best performances and light cyan cells correspond to the second best performances. All the NER based models outperform vanilla BERT for three out of four entity types. Overall, BERT-CRF performs the best.

### v4

@@ 6.2 Stage 2: Entity distillation and dictionary expansion
Package (PKG) pypy-configparser, account-plugin-twitter, gtkhtml3.2, pdfcrack, libfields-camlp4-dev-pf4q7 140062
Operating System (OS) SymbOS, Unix System III, NOS, Windows, Cosmic, kubuntu 877
Organization (ORG) launchpad, bugzilla, sourceforge, nokia, HP 379
Command (CMD) alias, arch, bzip, cat, clear 141
Error (ERR) No such process, No child processes, EFAULT, Bad address, EFBIG 124
Extension (EXT) .asm, .gz, .html, .log, .php 76
Peripheral (PRP) keyboard, mouse, printer, scanner, microphone 23
Software component (SOC) bios, driver, ui, ntfs, fat32 12
Architecture (ARC) x86, x64, 32-bit, 64-bit, amd64 7
@@
Human intervention: In this step, we avail human intervention to identify cases where a phrase has been incorrectly tagged as an entity. Active learning is particularly useful here, as the model can learn from the feedback provided by the human expert, improving its future predictions. We randomly sample a few automatically annotated bugs from all the bugs and check which phrases could potentially be non-entities. We encounter a challenge as the number of entities for categories like software components and peripherals is quite low in our entity list. To expand this dictionary, we utilize the software tool, TagMe [^17] https://sobigdata.d4science.org/web/tagme/tagme-help, to extract mentions from Wikipedia and consider them as entities. Along with that process, after receiving the annotated data from Stage 1 and Parts-of-speech filtered bugs (L) also with their IO mentions, a binary RoBERTa-based classifier is trained. We sample 100 bugs on a yearly basis, with our sample spanning from 2004 to 2019. Subsequently, we take up 100*16 = 1600 bugs from TagMe but not included in Stage 1 and feed them into our model for classification as either “entity” or “non-entity”. If the confidence for an entity being identified is 50% or greater, it undergoes manual labeling by a human to categorize it into one of the nine predefined entity types. Once the manual annotation is completed, the newly labeled data (d) is integrated into L, resulting in an updated dataset, (L $\cup$ d). The model is then retrained using this revised dataset, and the entire procedure is reiterated until the stopping criterion is met. The stopping criterion is defined as the absence of data in the list of entities identified by TagMe. Finally, our requirement for human intervention end up with around $\sim$3682 filtered mentions, a relatively small number considering the size of our dataset which contains 170K bugs (equivalent to $\sim$1.2 million entities). Distributions of identified entities are shown in Figure 3.
@@ 6.3 Stage 3: The NER model
We use a variety of NER models for obtaining the final tags. These include Linear-CRF [12], BilSTM-CRF [7], BERT-CRF [22], BERT-NER [17], RoBERTa-CRF [10], SpanBERT-CRF [20], and SoftNER [23]. All models are implemented to execute task-specific actions, with their performance evaluated using precision, recall, and F1-score. However, recall is highlighted in this paper as a representational metric for all classes in NER tasks. We train the existing CRF based models using our distantly labelled data obtained from Stage 1 and Stage 2. We note the different hyperparameters for the above models in Appendix. We also use LLMs in a zero-shot setting with instruction based prompt to extract the entities along with its start and end index from the paragraph (see section 10 for more details).
@@ 7 Heuristics
We use Part-of-Speech (POS) tagging as a heuristic to discard certain entities. In our process, we filter out the entities based on their POS tags. For example, we discard entities that are typically labeled as conjunctions, interjections, or prepositions as these are less likely to represent valid entities. This strategy ensures that only the most relevant words, such as nouns or proper nouns, are considered for entity recognition. This way, we can reduce noise in the data and improve the performance of our NER model. Furthermore, these POS tag-based heuristics help us in refining our entity list, leading to a more accurate and efficient distant supervision process in NER (See Table 3 for more samples).
8 Experimental setup
Training and test data preparation: In order to train the NER model in Stage 3, we need training data. For this purpose we select all the bug descriptions from the years 2004 through 2013 inspired by [6]. These descriptions, nearly 65K in number are automatically annotated using the first two stages of our framework. Thus, the training data only has auto-curated (aka silver) entity labels without any human involvement. For the test set we consider the data from the years 2016 to 2019. To assess the performance of the model, which has been trained using distantly labeled data, we perform human annotation of a subset of the test set. Human annotation: We present 500 bug descriptions to four domain experts, each claiming over three years of experience in the opensource ecosystem and package management. The resulting inter-annotator agreement among these annotators is 0.625 (see Appendix for more information). For the Launchpad QA dataset, we employ the same group of experts to annotate the entities in 500 question-answer pairs, ensuring consistency in our data annotation process across both datasets. Baselines: In addition to the NER models and zero-shot LLMs, we also use the combined output of Stage 1 and 2 as a baseline.
Metric for evaluation: To evaluate all the models, we compute the recall rate inspired by [26]. Recall rate is the proportion of actual entity types in ground truth that are accurately predicted. We use this metric rather than the F1 score for classwise evaluation in the main table in cognizance of the fact that human annotations are sometimes incomplete while the model is able to generate the correct entity type (see Section 9). Precisely, when dealing with software-related texts, overlooking an entity can lead to the omission of vital data. For instance, not recognizing an error code, a software package, or a particular function name can change the outcome from comprehending and addressing a problem to failing to do so. This makes capturing as many relevant entities as we can, which recall assesses, crucial. Further, software-related documents typically present crucial details that require exhaustive extraction for subsequent applications like bug tracking, requirement analysis, or code generation. Overlooking an entity in these contexts can have an adverse impact, underscoring the importance of recall as a key metric. Nevertheless, in order to make the results complete we also subsequently measure and report the recall and F1 score.
9 Results
@@
"This results in a serious compromise on the possibility of
running remote displays systems on ubuntu. In fact, the latter can only
@@
"SST will fail if donor has to send keyring.
Looks like the donor is trying to send the file
while so cat is still opening port $4$ on joiner…"
@@
We evaluate all the baselines for Ubuntu bug dataset, Launchpad QA, Fedora forum and Linux community dataset. We show classwise recall values for Ubuntu bug dataset and Launchpad QA in Table 4 while the overall performance in terms of macro-F1 score is shown in Table 5. Our experiments involve two different setups – (a) Human-Induced Training (HIn), and (b) Human-Only Training (HOn).
@@
Table 4: Recall rate for named entity recognition approaches. The first sub-row of each row shows the results for the Ubuntu bug dataset while the second sub-row shows the results for the Launchpad QA dataset (inference phase only). The best and the second best results in the first sub-row (bug dataset) are highlighted in dark and light green respectively. The best and the second best results in the second sub-row (QA dataset) are highlighted in dark and light purple respectively. BERT-CRF is significantly different from domain specific BERT-NER(except launchpad) and RoBERTa-CRF (*$p<0.05$). Table 5 compares methods for HIn based on other metrics.
In the case of a Human-Induced (HIn) configuration, we train the models using all the auto-annotated bug descriptions plus 10% (or $\sim$50 instances) human-annotated bug descriptions. These examples of human annotated data are incorporated into the training to provide the models with additional knowledge of gold annotations. For Human-Only training (HOn), only 10% of human annotations are used to train the models (see subsection 9.2). We do not include any auto-annotated data during HOn configuration training. In both setups, 10% of the human-annotated data is used for training, 20% for validation, and 70% for testing.
@@
Methods Pre-LLM era LLM era
Linear-CRF BiLSTM-CRF BERT-NER BERT-CRF RoBERTa-CRF spanBERT-CRF SoftNER GPT-3.5-Turbo GPT-4 Google BARD UniversalNER
Ubuntu
0.410 0.290 0.424 0.471 0.448 0.350 0.396 0.002 0.091 0.001 0.149
Launchpad
0.354 0.090 0.366 0.360 0.342 0.330 0.319 0.001 0.082 0.000 0.193
Fedora
0.403 0.323 0.429 0.495 0.417 0.213 0.314 0.009 0.018 0.003 0.191
Linux
0.441 0.285 0.449 0.507 0.477 0.302 0.371 0.009 0.033 0.003 0.204
Table 5: Comparison of methods for HIn based on macro-F1 Scores. BERT-CRF is significantly different from domain specific BERT-NER(except launchpad) and RoBERTa-CRF (*$p<0.05$). The best and the second best results are highlighted in dark and light purple respectively.
For evaluating overall results in terms of macro-F1 score we divide methods in two eras: the pre-LLM era and the LLM era. In the pre-LLM era, BERT-CRF consistently demonstrates superior performance across different datasets (Ubuntu (Bug), Launchpad (QA), Fedora (CQA), and Linux (CQA)), securing the highest macro-F1 scores highlighted in dark purple. Following closely, RoBERTa-CRF emerges as the second-best method in terms of performance, indicated by light purple highlights in the Ubuntu (Bug) and Linux (CQA) datasets. Transitioning to the LLM era, we notice a stark contrast in performance. Notably, the scores drastically drop, with GPT-3.5-Turbo, GPT-4, and Google BARD exhibiting significantly lower macro-F1 scores across all datasets, suggesting that despite their advanced capabilities, these models may not be directly optimized for the specific task of NER as compared to their predecessors in the pre-LLM era. However, UniversalNER demonstrates relatively better performance in this era, albeit still not reaching the effectiveness of the pre-LLM methods.
@@ 9.1 Progressive learning
Progressive learning [1] involves incrementally training machine learning models with increasing amounts of data. In our experiment, we divide the data into 25%, 50%, and 75% segments among equally distributed entities. Figure 4 illustrates the Recall rate distribution across different percentages of training data. We focus on the top three performing models and observe that BERT-CRF consistently improves with additional data, ultimately achieving the best performance when trained with 100% of the data.
@@ 10 Issues in finding distant labels using LLMs
Our research employs a uniform instruction-based approach to extract and tag entities from text, aiming to produce tagged entities with their indices in JSON format, as demonstrated in the Appendix with sample prompts. We assessed outputs from GPT-3.5, GPT-4, Google BARD, and UniversalNER, noting several key findings as follows. (a) Consistency and format: GPT-4 and Google BARD show consistent output patterns, unlike GPT-3.5-Turbo, which sometimes misses text or references lines inconsistently. GPT-3.5 and GPT-4 excel in entity accuracy but fall behind in index precision compared to BARD. (b) Invented entities: All models occasionally create new, untrained entity types in their outputs. (c) UniversalNER performance: Despite being specifically trained, UniversalNER struggles with entity accuracy. (d) Cost efficiency: Our distantly supervised approach offers a more cost-effective solution for entity tagging compared to direct model predictions. Some examples in appendix further illustrate these points.
@@ 11 Additional experiments for task-based evaluation
Relation extraction and NER are closely intertwined in many NLP tasks. They are often referred to as “sister" problems in the literature. Many studies, including [28, 33], explore these two tasks jointly, highlighting their interdependence. While relation extraction identifies and classifies the relationships between recognized entities, it often serves as a supplementary evaluation, building upon the foundational results of NER models.
@@ 11.1 Relation extraction
dependency), etc.), (b) conflict – conflict relation refers to a situation where two or more software components, packages, or entities cannot coexist or function harmoniously due to incompatibilities, overlapping functionalities, or conflicting configurations (e.g. triplets include (cups, Printers, conflicts), (Keyboard, Input Method Editor (IME), conflicts), etc.), affected version – these indicate which specific versions of a software are affected by reported issues or bugs (e.g. triplets include (Flatbed Scanner, macOS Mojave, affected version), (Error Code 134, sudo dpkg --configure -a, affected version), etc.), cause and effect – these relations refers to cases where one entity (cause) triggers another entity (effect) (e.g. triplets include (Error Code 502nginxcause and effect), (Error Code 401, openssh-server, cause and effect), etc.), and interaction/control – corresponds to relations where an entity exerts dynamic influence on / manipulates another entity (e.g. triplets include (apt, install, interaction/ control), (docker-ce, run, interaction/control), etc.). Dataset: For this experiment we use the earlier 500 bug description data that was manually annotated for the named entities. This data is further annotated with the relationships among the entities by 3 expert annotators with an inter-annotator agreement of 0.693. We end up with a total of 642 triplets, each consisting of a head entity, relationship, and tail entity. Out of these, 27% have a dependency relationship, 21% are of type affected versions while conflicts, cause and effect, and interaction/control types account for 4%, 17%, and 31% of the triplets respectively. Since the data for conflict type is very less, we ignore it for our experiments. Results: In Table 6 we show the results of the relation extraction task which is posed as a classification problem having the triplet (the head entity, the tail entity and the relation type) representation as the input and one of the relation classes (dependency, affected versions, cause and effect, and
Methods BERT-NER BERT-CRF RoBERTa-CRF
Ubuntu
0.119 0.130 0.101
Launchpad
0.073 0.072 0.098
Fedora
0.109 0.132 0.088
Linux
0.122 0.148 0.103
Table 7: Comparison of performance metrics for BERT-CRF, BERT-NER and RoBERTa-CRF methods across various categories using different datasets (Bug and Launchpad). The best and the second best results are highlighted in dark and light purple respectively.
interaction/control) as output. In the input, we draw the entity representations from our previously trained best NER models (i.e., trained BERT-CRF, trained BERT-NER, and trained RoBERTa-CRF) and compare their performance with vanilla BERT. Naturally, the rationale for employing the trained model as encoder is to acquire a more contextual representation than with the vanilla BERT. Further, we fine-tune the classifier layer with a very small amount of data (3%) to perform the relation classification. Dark cyan cells in Table 6 represent the best performances and light cyan cells correspond to the second best performances. All the NER based models outperform vanilla BERT for three out of four entity types. Overall, BERT-CRF performs the best.